# Barking Up The Syntactic Tree: Enhancing VLM Training with Syntactic Losses

### v1

@@ Abstract
Vision-Language Models (VLMs) achieved strong performance on a variety of tasks (e.g., image-text retrieval, visual question answering). However, most VLMs rely on coarse-grained image-caption pairs for alignment, relying on data volume to resolve ambiguities and ground linguistic concepts in images. The richer semantic and syntactic structure within text is largely overlooked. To address this, we propose HIerarchically STructured Learning (HIST) that enhances VLM training without any additional supervision, by hierarchically decomposing captions into the constituent Subject, Noun Phrases, and Composite Phrases. Entailment between these constituent components allows us to formulate additional regularization constraints on the VLM attention maps. Specifically, we introduce two novel loss functions: (1) Subject Loss, which aligns image content with the subject of corresponding phrase, acting as an entailment of standard contrastive/matching losses at the Phrase level; (2) Addition Loss, to balance attention across multiple objects. HIST is general, and can be applied to any VLM for which attention between vision and language can be computed; we illustrate its efficacy on BLIP and ALBEF. HIST outperforms baseline VLMs, achieving up to +9.8% improvement in visual grounding, +6.3% in multi-object referring segmentation, +1.1% in image-text retrieval, and +0.2% in visual question answering, underscoring the value of structuring learning in VLMs.
@@ 1 Introduction
Figure 1: Motivation. Existing VLM models train from image-caption pairs, requiring large corpora of data to establish word-region associations. The proposed HIerarchically STructured (HIST) learning framework, and associated losses, through syntactic analysis, decompose captions into hierarchy of noun phrases establishing entailment between them and the image and among the phrases themselves. Specifically, it extracts nouns (Subject level) from noun phrases and aligns them to the image, along with the corresponding noun phrases themselves (Noun Phrase level). In addition, it combines two noun phrases together into longer composite phrases and regularizes attention to be a sum of the constituent noun sub-phrases (Composite Combined Phrase level).
@@
Vision-Language Models (VLMs) [37, 27, 25, 26, 44, 2, 18, 23, 29, 33, 15], pretrained on vast amounts of image-text pairs, generalize effectively across a wide variety of downstream tasks, including image-text retrieval [31, 28], visual question answering [11], and visual grounding [17, 46]. However, VLMs primarily leverage a direct, coarse-grained alignment between image and caption pairs, often overlooking the deeper structures embedded in the textual data.
@@
Historically, the natural language processing (NLP) has made significant progress in understanding textual structure through the development of various parsing methods. Techniques such as syntactic parsing [30, 32], which decomposes sentences into hierarchical syntax trees, have been instrumental in identifying linguistic patterns and relationships between words [32]. These methods reveal the internal structure of text, allowing models to better capture the semantics within a sentence. Parsing has been effectively utilized in numerous tasks such as machine translation [3, 10] and sentence classification [20, 41], where understanding the hierarchical relationships in text provides substantial gains in accuracy. Syntactic parsing has also been shown to be useful in image-text retrieval [40], with order embeddings, and visual grounding [42]; preceding recent advances in LLMs and VLMs. However, such efforts have largely been abandoned more recently. Their efficacy and the general question of how richer textual representations can benefit VLM training remains open.
@@
To this end, we propose a novel HIerarchically STructured (HIST) learning framework that strengthens the connection between vision and language by embedding hierarchical text structures into the training process, without requiring annotation beyond image-text pairs. Our framework decomposes captions into noun phrases, each focusing on a distinct object within the image, and organizes these phrases into a three-level hierarchy: Subject Level, Noun Phrase Level, and Composite Combined Phrase Level, as depicted in Figure 1. The Combined Phrase level serves as the root, formed by joining two object-centric phrases using an “and” [^1] We do this, instead of using the raw structure of the caption, in order to get richer syntactic structure potentially derived from multiple captions.. The object-centric phrases themselves are represented at the (middle) Noun Phrase level. Finally, the parsed subject of each noun phrase form the Subject level and the leaf of the syntactic hierarchy.
@@
Entailment between these constituent components of the sentence and the image, allows us to formulate additional regularization constraints for training of VLMs. Specifically, we leverage three losses. At the Phrase Level we ensure alignment (entailment) between noun phrases and the image by leveraging standard VLM loses typically applied for image-caption pairs: image-text contrastive (ITC), image-text matching (ITM), and masked language modeling (MLM). At the Subject Level we similarly use ITC and ITM loses, but based on a single word noun. The motivation comes from the fact that a potentially much longer noun phrase, at the phrase level, may obfuscate text-image alignment by focusing more on adjectives (that maybe common among many phrases) as compared to the noun that tends to be more semantically important. The final, novel Addition loss, is formulated on emergent localization properties of VLMs [16, 24, 25] and requires the sum of the attention maps for the noun sub-phrases to be as close as possible to the combined composite phrase attention. This loss encourages the model to attend to multiple objects concurrently, rather than disproportionately focusing on the most prominent one. Collectively, these losses regularize training of VLMs, improving their performance, without requiring additional supervision beyond traditional image-caption pairs.
@@
  • Based on syntactic parsing, we propose a HIerarchically STructured weakly-supervised framework (HIST) that improved image-text alignment in VLM training.
  • We design two novel loss functions: (1) Subject Loss to ensure image alignment with both object-centric phrase and the subject, and (2) Addition Loss to encourage equal attention across multiple objects in a phrase.
Empirical results demonstrate the efficacy of our framework, showing an improvement of up to 9.8% on visual grounding and +6.3% on multi-object referring segmentation. Somewhat surprisingly, this approach, which was largely designed to improve spatial image-text alignment in VLMs, also benefits them more broadly with improvement of up to +1.1% on image-text retrieval, and +0.2% on visual question answering (VQA), underscoring the broader advantages of sentence structure modeling in VLMs.
Figure 2: The overall structure of HIST. We decompose image captions into object-centric phrases and build a three-level hierarchy – Subject level, Phrase level, and Composite Combined Phrase level. Entailment between these constituent components of the sentence and the image, allows us to formulate additional regularization constraints for training of VLMs. Specifically, we leverage three losses. At the Phrase Level we ensure alignment (entailment) between noun phrases and the image by leveraging standard VLM loses typically applied for image-caption pairs: image-text contrastive (ITC), image-text matching (ITM), and masked language modeling (MLM) or language modeling (LM), depending on the model. At the Subject Level we similarly use ITC and ITM loses, but focus on matching the image and the subject noun of the phrase. This prevents adjectives in the phrase from obscuring the text-image alignment. The final Addition loss requires the sum of the attention maps for the noun phrases to be as close as possible to the combined composite phrase attention. This loss encourages the model to attend to multiple objects concurrently, rather than disproportionately focusing on the most prominent one.
@@ 2 Related Work
Vision-Language Models (VLMs). Vision-language pre-training (VLP) has established itself as a cornerstone in multimodal learning, driving advancements across tasks. Various architectures have been developed to optimize these tasks, including dual-encoders like CLIP [33] and ALIGN [15] for scalable contrastive learning, fusion-encoders like ALBEF [24] and LXMERT [38] for early cross-modal integration, and unified transformers such as BLIP [25] and BEiT [4], which consolidate multiple capabilities into a single model with both encoder and decoder architecture. Recently, VLMs have also begun integrating large language models (LLMs), as seen in BLIP2 [26] and LLava [29], enhancing multimodal interactions by combining powerful language understanding with visual components. Key training objectives in VLP have converged around image-text contrastive learning [33, 24, 25, 23, 47, 37], image-text matching [24, 25, 44, 23, 47, 37, 18, 43], and (masked) language modeling [44, 23, 47, 37, 6, 23, 2, 25], enabling refined and versatile multimodal understanding.
@@
Despite these strides, most VLMs rely on pairwise image-text matching, often overlooking hierarchical structures in data, such as relationships among sub-phrases. Addressing this, our approach introduces hierarchical alignment, regularizing VLM learning with these nuanced relationships. This method enhances the model’s ability to localize and ground language-described entities in images; without the need for additional annotations or data.
@@
Vision-Language Learning with Language Structure Supervised Losses. Several prior works have explored the benefits of leveraging language structure to enhance vision-language understanding. Early studies, such as [40], demonstrated the advantage of encoding textual hierarchical relationships (like generalization and entailment) in embedding spaces to improve tasks such as hypernym prediction, image-caption retrieval, and textual entailment. Another work [42] applied compositionality and exclusivity constraints, across noun phrases, to attention maps to improve visual grounding. However, these works were limited to smaller models and datasets, leaving open the question of to what extent these findings hold for the recent VLMs.
@@
More recently, [8] introduces a semi-supervised approach that aligns image regions with corresponding noun phrases in captions, improving vision-language models (VLMs) on tasks such as visual question answering, visual commonsense reasoning, and visual grounding. Related, [5] examines how attributes, such as color, size, and texture, within image captions influence the learning and performance of vision-language models on open-vocabulary object detection and fine-grained text-region retrieval. SelfEQ [14] fine-tunes VLMs to produce consistent visual explanations across paraphrased captions, showing marked improvements in visual grounding tasks.
Conceptually related [22] and [48] apply constraints to GradCAM maps according to linguistic structure, with the aim of enhancing visual grounding. In [48], dependency parsing on image captions is used to identify distinct objects and prevents overlapping attention maps between them, thereby improving model performance on visual grounding. [22] designs specific losses for referring segmentation, enforcing that words within the same noun phrase have similar GradCAM maps while suppressing attention on incorrect object regions based on relationships with neighboring words (building on observations in [42]).
@@
We are inspired by [14] and [22], however, propose a formulation that models complementary sentence structure in terms of hierarchy of phrase and sub-phrase relations. As such, we posit that our loses and objectives can be combined with consistency [14] and exclusion [22] loses designed in those works. Moreover, we are the first, to our knowledge, to show that such regularization losses not only improve visual localization, but general VLM understanding.
@@ 3 Method
Our approach aims to improve the visual-textual alignment of VLMs by fine-tuning them with a novel HIerarchicaly STructured (HIST) set of losses for the constituent text phrases. Specifically, we decompose image captions into object-centric phrases and build a three-level hierarchy – Subject level, Phrase level, and Composite Combined Phrase level – as shown in Figure 1; with the steps specified in Algorithm 1 and Section 3.2. The hierarchical structure allows us to design training objectives that achieve more granular alignment between visual and textual features.
@@
In Section 3.1, we first introduce preliminaries and the base model architectures to which we apply proposed HIST. These base models are typically optimized with three widely used losses formulated atop of image-text pairs: Image-Text Contrastive (ITC), Image-Text Matching (ITM), and Masked Language Modeling (MLM). We use these losses in conjunction with our hierarchically structured losses to fine-tune the base models. In Section 3.2, we present our hierarchical alignment objectives: the Subject loss and the Addition loss. The Subject loss aims to align the main subject of each noun phrase with its corresponding visual representation, enforcing consistency between subjects and corresponding phrases. The Addition loss further guides the model to attend to multiple objects within the same image by aggregating Grad-CAM maps across different phrases, ensuring that the model’s attention covers all relevant objects. This hierarchical loss structure allows our model to align fine-grained visual and textual details while maintaining global coherence across the image-text pair.
@@
Experimentally, we show that these additional losses, that require no added supervision beyond image-text pairs, lead to significant improvement on grounding. Further, we show that they moderately improve performance on a variety of other downstream VLM tasks (see Sec. 4.3).

### v2

# Barking Up The Syntactic Tree: Enhancing Emergent Visual Grounding with Syntactic Losses

@@ Abstract
Vision-Language Models (VLMs) implicitly learn to associate image regions with words from large-scale training data, demonstrating an emergent capability for grounding concepts without dense annotations [14, 51, 18]. However, the coarse-grained supervision from image-caption pairs is often insufficient to resolve ambiguities in object-concept correspondence, even with enormous data volume. Rich semantic and syntactic structures within the text modality have been overlooked as sources of supervision. Starting from contrastive architectures (BLIP and ALBEF) that show strong intrinsic grounding abilities, we propose HIerarchically STructured Learning (HIST). HIST enhances spatial vision-language alignment without using additional human annotations, by hierarchically decomposing captions into the constituent Subjects, Phrases, and Composite Phrases, and enforcing entailment relation between a parent and its children in the hierarchy. Specifically, we introduce two novel loss functions: (1) Subject Loss, which aligns image content with the subject of the corresponding phrase, acting as an entailment of standard contrastive/matching losses at the Phrase level; (2) Composition Loss, to balance attention across multiple objects. HIST is general, and can be applied to any VLM for which attention between vision and language can be computed. Compared to baseline VLMs, HIST achieves up to +9.8% improvement in visual grounding and +6.3% in multi-object referring segmentation. Surprisingly, the improved spatial grounding leads to improvements in other downstream VLM tasks: +1.1% in image-text retrieval, and +0.2% in visual question answering.
@@ 1 Introduction
Figure 1: Motivation. Existing VLM models train from unstructured image-caption pairs. The proposed HIerarchically STructured (HIST) learning framework decompose captions into hierarchy of phrases, establishing entailment between phrases and the image, and among the phrases themselves using proposed losses. Specifically, it extracts subjects (Subject level) from phrases and aligns them to the image, along with the corresponding phrases themselves (Phrase level). In addition, it combines two phrases together into longer composite phrase and regularizes attention to be a sum of the constituent sub-phrases (Composite Phrase level).
@@
Vision-Language Models (VLMs) [45, 30, 28, 29, 55, 2, 21, 26, 33, 39, 17], pretrained on vast amounts of image-text pairs, generalize effectively across a wide variety of downstream tasks, including image-text retrieval [36, 32], visual question answering [12], and visual grounding [20, 58]. These models demonstrate emergent SoTA grounding abilities, even though they are not trained directly to perform grounding or segmentation. However, VLMs primarily leverage a direct, coarse-grained alignment between image and caption pairs when training, often overlooking the deeper structures embedded in the textual data.
@@
Historically, the natural language processing (NLP) has made significant progress in understanding textual structure through the development of various parsing methods. Techniques such as syntactic parsing [35, 37], which decomposes sentences into hierarchical syntax trees, have been instrumental in identifying linguistic patterns and relationships between words [37]. These methods reveal the internal structure of text, allowing models to better capture the semantics within a sentence. Parsing has been effectively utilized in numerous tasks such as machine translation [3, 11] and sentence classification [23, 52], where understanding the hierarchical relationships in text provides substantial gains in accuracy. Preceding recent advances in LLMs and VLMs, syntactic parsing has also been shown to be useful in image-text retrieval [50], with order embeddings, and visual grounding [53]. However, the general question of how richer textual representations can benefit spatial image-text alignment, or implicit grounding, in current large-scale VLMs training remains open.
@@
To this end, we propose a novel HIerarchically STructured (HIST) framework incorporates hierarchical text structures into training, without requiring annotation beyond image-text pairs. HIST decomposes captions into phrases, each focusing on a distinct object within the image, and organizes these phrases into a three-level hierarchy: Subject Level, Phrase Level, and Composite Phrase Level, as depicted in Figure 1. The Composite Phrase level serves as the root, formed by joining two phrases using an “and” [^1] We do this, instead of using the raw structure of the caption, in order to get richer syntactic structure potentially derived from multiple captions.. The phrases themselves are represented at the (middle) Phrase level. Finally, the parsed subject of each phrase form the Subject level and the leaf of syntactic hierarchy.
@@
We formulate regularization constraints based on entailment relations between these components. Specifically, we leverage three losses. At the Phrase Level we ensure alignment (entailment) between phrases and the image by leveraging standard VLM loses typically applied for image-caption pairs. At the Subject Level we use similar losses, but based on a single word. The motivation comes from the fact that a potentially much longer phrase, at the phrase level, may obfuscate text-image alignment by focusing more on adjectives (that maybe common among many phrases) as compared to the noun or actor that tends to be more semantically important. The final, novel Composition loss, is formulated on emergent localization properties of VLMs [18, 27, 28] and requires the sum of the attention maps for the sub-phrases to be as close as possible to the composite phrase attention. This loss encourages the model to attend to multiple objects concurrently, rather than disproportionately focusing on the most prominent one. Collectively, these losses regularize the training of VLMs, improving their performance on visual grounding, without requiring additional supervision beyond traditional image-caption pairs.
Empirically, HIST obtains improvements of up to 9.8% on general visual grounding and +6.3% on multi-object referring segmentation. Somewhat surprisingly, this approach also benefits models more broadly with improvement of up to +1.1% on image-text retrieval, and +0.2% on visual question answering (VQA), underscoring the broader advantages of sentence structure modeling in VLM training.
@@
  • Leveraging syntactic parsing, we propose a HIerarchically STructured weakly-supervised (HIST) framework that improves spatial image-text alignment in VLMs.
  • We design two novel loss functions: (1) Subject Loss to ensure image alignment with both object-centric phrase and the subject, and (2) Composition Loss to encourage equal attention across multiple objects in a phrase.
We will release all models, data and code upon acceptance of the paper.
Figure 2: The overall structure of HIST. We decompose image captions into object-centric phrases and build a three-level hierarchy – Subject level, Phrase level, and Composite Phrase level. Entailment between these constituent components of the sentence and the image, allows us to formulate additional regularization constraints for training of VLMs. Specifically, we leverage three losses. At the Phrase Level we ensure alignment (entailment) between phrases and the image by leveraging standard VLM loses typically applied for image-caption pairs: image-text contrastive (ITC), image-text matching (ITM), and masked language modeling (MLM) or language modeling (LM), depending on the model. At the Subject Level we similarly use ITC and ITM loses (SITC and SITM), but focus on matching the image and the subject of the phrase. This prevents adjectives in the phrase from obscuring the text-image alignment. The final Composition loss (LCOMP) requires the sum of the attention maps for the phrases to be as close as possible to the composite phrase attention. This loss encourages the model to attend to multiple objects concurrently, rather than disproportionately focusing on the most prominent one. In computing this loss we leverage a product of cross-attention map and GradCam illustrated for each layer and head.
@@ 2 Related Work
Vision-Language Models (VLMs). Vision-language pre-training has established itself as a cornerstone in multimodal learning, driving advancements across tasks. Various architectures have been developed to optimize these tasks, including dual-encoders like CLIP [39] and ALIGN [17] for scalable contrastive learning, fusion-encoders like ALBEF [27] and LXMERT [48] for early cross-modal integration, and unified transformers such as BLIP [28] and BEiT [4], which consolidate multiple capabilities into a single model with both encoder and decoder architecture. Recently, VLMs have also begun integrating large language models (LLMs), as seen in BLIP2 [29] and LLava [33], enhancing multimodal interactions by combining powerful language understanding with visual components. Key pretraining objectives have converged around image-text contrastive learning [39, 27, 28, 26, 59, 45], image-text matching [27, 28, 55, 26, 59, 45, 21, 54], and (masked) language modeling [55, 26, 59, 45, 7, 26, 2, 28], enabling refined and versatile multimodal understanding.
@@
Despite advances, most VLMs rely on pairwise image-text matching, overlooking hierarchical structures in data, such as relationships among sub-phrases. Addressing this, our approach introduces hierarchical alignment, regularizing VLM learning with these nuanced relationships. This method enhances the model’s ability to localize and ground language-described entities in images; without the need for additional annotations or data.
@@
Learning with Language Structure Supervised Losses. Several prior works have explored the benefits of leveraging language structure to enhance vision-language understanding. Early studies, such as [50], demonstrated the advantage of encoding textual hierarchical relationships (like generalization and entailment) in embedding spaces to improve tasks such as hypernym prediction, image-caption retrieval, and textual entailment. Alternatively, [53] applied compositionality and exclusivity constraints across phrases to attention maps to improve visual grounding. However, these works were limited to smaller models and datasets, leaving open the question of extent these findings hold for VLMs.
@@
More recently, [9] introduces a semi-supervised approach that aligns image regions with corresponding phrases in captions, improving VLMs on tasks such as visual question answering, visual commonsense reasoning, and visual grounding; [5] examines how attributes, such as color, size, and texture, within image captions influence the learning and performance on open-vocabulary object detection and fine-grained text-region retrieval. SelfEQ [16] finetunes VLMs to produce consistent visual explanations across paraphrased captions, showing marked improvements in visual grounding tasks. Conceptually related, [25] and [60] apply linguistic constraints to GradCAM maps to enhance visual grounding. [60] identifies distinct objects from dependency parsing of image captions and prevents their attention maps from overlapping. Building off [53], [25] enforces that words within the same phrase have similar GradCAM maps while suppressing attention on incorrect object regions for referring image segmentation.
@@
We are inspired by [16] and [25], however, propose a formulation that models complementary sentence structure in terms of hierarchy of phrase and sub-phrase relations. As such, we posit that our losses and objectives can be combined with consistency [16] and exclusion [25] loses designed in those works; but such exploration is beyond the scope of this work. Moreover, we are the first, to our knowledge, to show that such regularization losses not only improve visual localization, but general VLM understanding.
@@ 3 Method
Our approach aims to improve the spatial image-text alignment of VLMs by fine-tuning them with a novel HIerarchicaly STructured (HIST) set of losses for the constituent text phrases. Specifically, we decompose image captions into object-centric phrases and build a three-level hierarchy – Subject level, Phrase level, and Composite Phrase level – as shown in Figure 1; with the steps specified in Algorithm 1 and Section 3.3. The hierarchical structure allows us to design training objectives that achieve more granular alignment between visual and textual features.
@@
In Section 3.1, we first discuss the performance of common classes of VLMs on weakly supervised visual grounding and our selection of backbone models. Then in Section 3.2, we introduce preliminaries and model architectures of our backbone models. These base models are typically optimized with three widely used losses formulated atop of image-text pairs: Image-Text Contrastive (ITC), Image-Text Matching (ITM), and Masked Language Modeling (MLM) or Language Modeling (LM). We use these losses in conjunction with our hierarchically structured losses to fine-tune the base models. In Section 3.3, we present our hierarchical alignment objectives: the Subject loss and the Composition loss. The Subject loss aims to align the main subject of each phrase with its corresponding visual representation, enforcing consistency between subjects and corresponding phrases. The Composition loss further guides the model to attend to multiple objects within the same image by aggregating Grad-CAM maps across different phrases, ensuring that the model’s attention covers all relevant objects. This hierarchical loss structure allows our model to align fine-grained visual and textual details while maintaining global coherence across the image-text pair.
@@
Experimentally, we show that these additional losses, that require no added supervision beyond image-text pairs, lead to significant improvement on grounding. Further, we show that they moderately improve performance on a variety of other downstream VLM tasks (see Section 4.3).
3.1 On the Choice of VLMs
Our formulation, HIST, is applicable to any VLM model that leverages spatial image-text attention. However, we focus on experimenting on state-of-the-art (SoTA) models. In the task of weakly supervised visual grounding, the SoTA models are typically trained on contrastive losses, including ALBEF [27], CLIP [39], and BLIP [28]. This also allows us to fairly compare to closely related work (e.g., [16]).
More recent VLMs, such as LLaVA [33] and Qwen [38], are primarily trained on the language modeling loss. Though they achieve high performance on image captioning and visual question answering, they are weaker in the task of weakly-supervised visual grounding than BLIP and ALBEF (see summary of results in Table 1). Recent works [19] also illustrate that LLaVA like variants suffer from attention sinks – high attention values in non-informative regions - which inversely affect the grounding accuracy when utiling attention maps. Still, to illustrate broader generality of our ideas, we also evaluate HIST on a LLaVA variant, TinyLLaVA [65], in Section 5.
RefCOCO+
Method Size Training val TestA TestB
Model trained with Contrastive Losses
[…]
Model trained with the Language Modeling Loss
[…]
Table 1: Zero-shot Performance of VLMs on Weakly-supervised Visual Grounding. Performance (IoU@0.5) of VLMs with contrastive losses (top) and language modeling losses (bottom) on the referring expression comprehension. Contrastive models like ALBEF, while smaller, exhibit much better performance.
3.2 Base Models and Objectives
ALBEF [27] and BLIP [28] both consist of an unimodal image encoder $\boldsymbol{\phi_{v}}$, a unimodal text encoder $\boldsymbol{\phi_{t}}$, and a multimodal encoder $\boldsymbol{\phi_{m}}$. Both models utilize the three commonly used objectives for learning visual and textual representations: image-text contrastive (ITC), image-text matching (ITM), masked language modeling (MLM) or language modeling (LM). We provide a brief overview of these objectives here as we also leverage them during our training.
Image-Text Contrastive Learning (ITC). The ITC objective $\mathcal{L}_{\text{itc}}$ aims to align unimodal representations by maximizing similarity of visual and textual representations for matching image-text pairs and minimizing it for non-matching pairs. Following MoCo [15], a momentum encoder is introduced to produce stable features representation queues and create soft labels to account for the potential positives in the negative pairs.
Image-Text Matching (ITM). The ITM objective $\mathcal{L}_{\text{itm}}$ predicts whether an image-text pair $(\mathbf{I},\mathbf{T})$ is positive (matched) or negative (not matched). This is achieved by passing the fused [CLS] embedding from the multimodal encoder through a fully connected layer, followed by a softmax function to predict the binary probability $p^{\text{itm}}(\mathbf{I},\mathbf{T})$. The ITM loss minimizes the cross-entropy between ground-truth label (matched or not) and the predicted probability.
Masked Language Modeling (MLM) or Language Modeling (LM). ALBEF [27] leverages Masked Language Modeling (MLM), $\mathcal{L}_{\text{mlm}}$ which integrates contextual information from both the text and the paired image to predict masked tokens within the text. BLIP [28], on the other hand, use a language modeling loss $\mathcal{L}_{\text{lm}}$ focused on predicting the next word token rather than randomly masking words within sentences. Essentially, this approach uses a cross-entropy loss for next-word /masked token prediction.